OpenStudy (anonymous):

Prove That:- tan(π/12) * tan(5π/12) * tan(7π/12) * tan(11π/12) = 1

OpenStudy (anonymous):

\[\tan(\frac{11 \pi}{12}) = \tan(\pi - \frac{ \pi}{12}) = -\tan(\frac{\pi}{12})\]

OpenStudy (anonymous):

\[\tan(\frac{7 \pi}{12}) = \tan(\pi - \frac{ 5 \pi}{12}) = -\tan(\frac{5 \pi}{12})\]

OpenStudy (anonymous):

yeah, i've understood dat portion...but could you plz solve the whole problem...i mean i'm nt getting it completely

OpenStudy (anonymous):

Where are you not getting it?

OpenStudy (anonymous):

\[LHS: [\tan(\frac{\pi}{12})\tan(\frac{5 \pi}{12})]^2\]

OpenStudy (anonymous):

yeah, then? how come the answer is 1?

OpenStudy (anonymous):

\[\tan(\frac{5 \pi}{12}) = \tan(\frac{\pi}{2} - \frac{\pi}{12}) = \cot(\frac{\pi}{12})\]

OpenStudy (anonymous):

ooooooooo yeah, i got it :P

OpenStudy (anonymous):

\[LHS: [\tan(\frac{\pi}{12})\cot(\frac{ \pi}{12})]^2\]

OpenStudy (anonymous):

Rest you know..

OpenStudy (anonymous):

thank you, very much! :)

OpenStudy (anonymous):

You are welcome dear.. :)